# Quasiprobability Thermodynamic Uncertainty Relation

### v1

@@ Abstract
Example.— We illustrate the above discussion through the model used in Ref. [54]. Its Hamiltonian $H$ has two $N$-fold degenerate eigenvalues, 0 and $\omega$. The jump operators represent simultaneous jumps between energy levels; with an eigenbasis $\{|g,j\rangle,|e,j\rangle\}_{j=1}^{N}$ such that $H|g,j\rangle=0$ and $H|e,j\rangle=\omega|e,j\rangle$, they are defined as $L_{+}=\sqrt{\gamma_{+}}\sum_{j,j^{\prime}}|e,j\rangle\langle g,j^{\prime}|$ and $L_{-}=\sqrt{\gamma_{-}}\sum_{j,j^{\prime}}|g,j\rangle\langle e,j^{\prime}|$. Anomalous scaling $m_{H}(\rho)=O(N^{2})$ is observed if we set $\rho=\rho^{+}=p_{g}|g,+\rangle\langle g,+|+p_{e}|e,+\rangle\langle e,+|$ with $|s,+\rangle=\frac{1}{\sqrt{N}}\sum_{j}|s,j\rangle$ for $s=e,g$ (see Supplemental Material [74]). The state $\rho^{+}$ has a large $l_{1}$-coherence $C_{l_{1}}(\rho^{+})=O(N)$, defined as $C_{l_{1}}(\rho)=\sum_{s,j,s^{\prime},j^{\prime}|(s,j)\neq(s^{\prime},j^{\prime})}|\langle s,j|\rho|s^{\prime},j^{\prime}\rangle|$, to which Ref. [54] attributed the dissipationless current. Now, note that the basis $\{|g,j\rangle,|e,j\rangle\}_{j=1}^{N}$, defining the $l_{1}$-coherence, is classical since there are only two jumps regardless of $N$ and, for example, $\langle e,j^{\prime}|L_{+}|g,j\rangle=\sqrt{\gamma_{+}}=O(1)$. On the other hand, if we consider a non-classical basis including $|g,+\rangle$ and $|e,+\rangle$ (note that $\langle e,+|L_{+}|g,+\rangle=\sqrt{\gamma_{+}}N$), the state $\rho^{+}$ has no coherence regarding this basis. This fact shows that $m_{H}(\rho^{+})$ may exhibit anomalous scaling (that is, we may have dissipationless current) even if $\rho^{+}$ has no coherence if the referred basis is not classical. Nonetheless, we see that $\rho^{+}$ fulfills either (Q1) or (Q2) for any eigenbasis, from the general discussion.
@@
In addition, we can construct a state $\rho^{-}$ that has as much coherence regarding the classical eigenbasis as $\rho^{+}$ but does not yield dissipationless current, violating both of the quantumness conditions, (Q1) and (Q2). Such a state is constructed by replacing $|s,+\rangle$ in the definition of $\rho^{+}$ with $|s,-\rangle=\frac{1}{\sqrt{N}}\sum_{j}(-1)^{j}|s,j\rangle$. While it has the same amount of $l_{1}$-coherence as $\rho^{+}$, the integrated fluxes and average escape rate read $\mathcal{T}_{s^{\prime}s}(\rho^{-})=\gamma_{s}p_{s}\chi(N)$ for $s\neq s^{\prime}$, $\mathcal{T}_{ss}(\rho^{-})=\frac{\gamma_{s}p_{s}}{2}(N-\chi(N))$, and $\bar{\lambda}(\rho^{-})=\frac{\gamma_{+}p_{g}+\gamma_{-}p_{e}}{2}(N+\chi(N))$, where $\gamma_{g}=\gamma_{+}$, $\gamma_{e}=\gamma_{-}$, and $\chi(N)$ is the remainder when $N$ is divided by $2$ [74]. Therefore, $\rho^{-}$ does not satisfy either (Q1) and (Q2), and leads to $m_{H}(\rho^{-})=O(1)$. Hence, anomalous dissipationless current does not occur for $\rho^{-}$, which cannot be understood by looking at the coherence alone.
@@
Discussion.— We have explored fundamental connections between quantum thermodynamics and quasiprobability; in particular, we have derived a quantum TUR evaluating fluctuations through the TMH quasiprobability. The TUR is valid for any state $\rho$ and observable $X$, as long as the dynamics satisfy a few assumptions necessary to discuss thermodynamics. As discussed, our TUR is complementary to existing TURs, which adopt different approaches when quantifying fluctuations. We have also demonstrated that when discussing anomalous scaling of fluctuations in the TUR, which brings reduction of dissipation beyond classical limitations, non-classical behaviors of quasiprobability are more essential than quantum coherence.
@@
To the best of our knowledge, our study is the first to bridge quasiprobabilities and the thermodynamic trade-off relations, recently discovered in stochastic thermodynamics [75]. In addition to the TUR, several fundamental relations have been unveiled in the field, such as thermodynamic speed limits [76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86], bounds on asymmetry [87, 88, 89, 90], and fluctuation-response inequalities [91, 92, 93]. We expect that quasiprobabilities will help explore non-classical features in their quantum extensions.
@@ Acknowledgements.
Acknowledgments.— K.Y. is supported by the Special Postdoctoral Researchers Program at RIKEN. R.H. is supported by JSPS KAKENHI Grant No. JP24K16982. K.Y. and R.H. are supported by JST ERATO Grant No. JPMJER2302, Japan.
@@
$\mathbb{J}_{k}(\rho)$ $\coloneqq\begin{pmatrix}O&J_{-k}(\rho)\\ J_{k}(\rho)&O\end{pmatrix}$, (35)
$\mathbb{F}_{k}(\rho)$ $\coloneqq\begin{pmatrix}O&F_{-k}(\rho)\\ F_{k}(\rho)&O\end{pmatrix}$, (36)
@@
$J_{k}(\rho)$ $\coloneqq\frac{1}{2}(\gamma_{k}\tilde{L}_{k}\rho-\gamma_{-k}\rho\tilde{L}_{k}),$ (37)
$F_{k}(\rho)$ $\coloneqq s_{k}\tilde{L}_{k}+[\tilde{L}_{k},\ln\rho].$ (38)
@@
$\mathbb{J}(\rho)\coloneqq\bigoplus_{k}\mathbb{J}_{k}(\rho),\quad\mathbb{F}(\rho)\coloneqq\bigoplus_{k}\mathbb{F}_{k}(\rho),$ (39)
@@
$\dot{\Sigma}(\rho)=\langle\mathbb{J}(\rho),\mathbb{F}(\rho)\rangle,$ (40)
@@
$\nabla_{\mathbb{L}}A\coloneqq[I_{\mathfrak{h}}\otimes A,\mathbb{L}],$ (41)
@@
$\mathbb{L}\coloneqq\bigoplus_{k}\begin{pmatrix}O&\tilde{L}_{-k}\\ \tilde{L}_{k}&O\end{pmatrix}$. (42)
@@
$\Gamma=\bigoplus_{k}\begin{pmatrix}\gamma_{k}/2&0\\ 0&\gamma_{-k}/2\end{pmatrix}$. (44)
@@
$\dot{\Sigma}(\rho)=\lVert\mathbb{F}(\rho)\rVert_{\Gamma\otimes\rho}^{2}.$ (45)
@@
$\dot{\Sigma}(\rho)\lVert\nabla_{\mathbb{L}}X\rVert_{\Gamma\otimes\rho}^{2}\geq|\langle\mathbb{F}(\rho),\nabla_{\mathbb{L}}X\rangle_{\Gamma\otimes\rho}|^{2}.$ (46)
@@
Appendix A Computational details of Example
We provide the details of the computation in the example in the main text. We aim to compute $\mathcal{T}_{s^{\prime}s}(\rho)$ and $\bar{\lambda}(\rho)$ for $\rho=\rho^{\pm}$ and show $m_{H}(\rho^{+})=O(N^{2})$ and $m_{H}(\rho^{-})=O(1)$. Here, the model we consider consists of Hamiltonian $H=\omega\sum_{j}|e,j\rangle\langle e,j|$ and jump operators $L_{+}=\sqrt{\gamma_{+}}\sum_{j,j^{\prime}}|e,j\rangle\langle g,j^{\prime}|$ and $L_{-}=\sqrt{\gamma_{-}}\sum_{j,j^{\prime}}|g,j\rangle\langle e,j^{\prime}|$. The states of interest $\rho^{\pm}$ are generated as $\rho^{\pm}=\sum_{s=g,e}p_{s}|s,\pm\rangle\langle s,\pm|$, where $|s,\pm\rangle=\frac{1}{\sqrt{N}}\sum_{j}(\pm 1)^{j}|s,j\rangle$ and $p_{s}$ satisfies $p_{s}\geq 0$ and $p_{g}+p_{e}=1$.
@@ Appendix A Computational details of Example
$T_{s^{\prime}j^{\prime}sj}(\rho)=\frac{1}{2}\mathop{\mathrm{tr}}\nolimits\big{(}\{\mathcal{D}^{\dagger}\Pi_{s^{\prime}j^{\prime}},\Pi_{sj}\}\rho\big{)}.$ (47)
@@
$\sigma_{s,j,j^{\prime}}$ $\coloneqq\frac{1}{2}\big{(}|s,j\rangle\langle s,j^{\prime}|+|s,j^{\prime}\rangle\langle s,j|\big{)},$ (48)
$\Lambda_{s,j}$ $\coloneqq\sum_{j^{\prime}}\sigma_{s,j,j^{\prime}}.$ (49)
@@
$\sigma_{s,j,j^{\prime}}=\sigma_{s,j^{\prime},j},\;\;\sigma_{s,j,j}=\Pi_{s,j},$ (50)
$\mathop{\mathrm{tr}}\nolimits(\sigma_{s,j,j^{\prime}}\rho)=\mathrm{Re}\langle s,j|\rho|s,j^{\prime}\rangle,$ (51)
$\sum_{j}\Lambda_{s,j}=\sum_{j,j^{\prime}}|s,j\rangle\langle s,j^{\prime}|,$ (52)
$\frac{1}{2}\sum_{j^{\prime}}\{\Lambda_{s^{\prime},j^{\prime}},\Pi_{s,j}\}=\delta_{s,s^{\prime}}\Lambda_{s,j}.$ (53)
@@
$L_{+}^{\dagger}\Pi_{s,j}L_{+}$ $=\gamma_{+}\delta_{s,e}\sum_{j^{\prime}}\Lambda_{g,j^{\prime}},$ (54)
$L_{-}^{\dagger}\Pi_{s,j}L_{-}$ $=\gamma_{-}\delta_{s,g}\sum_{j^{\prime}}\Lambda_{e,j^{\prime}},$ (55)
@@
$L_{+}^{\dagger}L_{+}$ $=\gamma_{+}N\sum_{j^{\prime}}\Lambda_{g,j^{\prime}},$ (56)
$L_{-}^{\dagger}L_{-}$ $=\gamma_{-}N\sum_{j^{\prime}}\Lambda_{e,j^{\prime}}.$ (57)
@@
$+\delta_{s,g}\bigg{(}\gamma_{-}\sum_{j^{\prime}}\Lambda_{e,j^{\prime}}-\gamma_{+}N\Lambda_{g,j}\bigg{)}.$ (58)
@@
$+\delta_{s^{\prime},g}\delta_{s,e}\gamma_{-}\Lambda_{e,j}-\frac{\delta_{s^{\prime},g}\delta_{s,g}}{2}\gamma_{+}N(\delta_{j,j^{\prime}}\Lambda_{g,j}+\sigma_{g,j,j^{\prime}}).$ (59)
@@
$T_{ej^{\prime}gj}(\rho)$ $=\gamma_{+}\sum_{j^{\prime\prime}}\mathrm{Re}\langle g,j|\rho|g,j^{\prime\prime}\rangle$ (60)
$T_{gj^{\prime}ej}(\rho)$ $=\gamma_{-}\sum_{j^{\prime\prime}}\mathrm{Re}\langle e,j|\rho|e,j^{\prime\prime}\rangle$ (61)
@@
$\phantom{===}+\delta_{j,j^{\prime}}\sum_{j^{\prime\prime}}\mathrm{Re}\langle g,j|\rho|g,j^{\prime\prime}\rangle\bigg{)}$ (62)
@@
$\phantom{===}+\delta_{j,j^{\prime}}\sum_{j^{\prime\prime}}\mathrm{Re}\langle e,j|\rho|e,j^{\prime\prime}\rangle\bigg{)}.$ (63)
@@
$T_{ej^{\prime}gj}(\rho^{+})=p_{g}\gamma_{+},\;\;T_{ej^{\prime}gj}(\rho^{-})=(-1)^{j+1}p_{g}\gamma_{+}\frac{\chi(N)}{N},$ (64)
@@
$\mathcal{T}_{eg}(\rho^{+})=p_{g}\gamma_{+}N^{2},\quad\mathcal{T}_{eg}(\rho^{-})=p_{g}\gamma_{+}\chi(N).$ (65)
@@
$T_{gj^{\prime}gj}(\rho^{+})$ $=-\frac{\gamma_{+}p_{g}}{2}(1+\delta_{j,j^{\prime}}N)$ (66)
$T_{gj^{\prime}gj}(\rho^{-})$ $=-\frac{(-1)^{j}\gamma_{+}p_{g}}{2}((-1)^{j^{\prime}}-\delta_{j,j^{\prime}}\chi(N)).$ (67)
@@
$\mathcal{T}_{gg}(\rho^{+})$ $=-\frac{\gamma_{+}p_{g}}{2}N(N-1),$ (68)
$\mathcal{T}_{gg}(\rho^{-})$ $=\frac{\gamma_{+}p_{g}}{2}(N-\chi(N))$ (69)
@@
$\sum_{j,j^{\prime}(\neq j)}(-1)^{j}(-1)^{j^{\prime}}$ $=\bigg{(}\sum_{j}(-1)^{j}\bigg{)}^{2}-\sum_{j}(-1)^{2j}$ (70)
$=\chi(N)-N.$ (71)

### v2

@@ Abstract
Example.— We illustrate the above discussion through the model used in Ref. [57]. Its Hamiltonian $H$ has two $N$-fold degenerate eigenvalues, 0 and $\omega$. The jump operators represent simultaneous jumps between energy levels; with an eigenbasis $\{|g,j\rangle,|e,j\rangle\}_{j=1}^{N}$ such that $H|g,j\rangle=0$ and $H|e,j\rangle=\omega|e,j\rangle$, they are defined as $L_{+}=\sqrt{\gamma_{+}}\sum_{j,j^{\prime}}|e,j\rangle\langle g,j^{\prime}|$ and $L_{-}=\sqrt{\gamma_{-}}\sum_{j,j^{\prime}}|g,j\rangle\langle e,j^{\prime}|$. Anomalous scaling $m_{H}(\rho)=O(N^{2})$ is observed if we set $\rho=\rho^{+}=p_{g}|g,+\rangle\langle g,+|+p_{e}|e,+\rangle\langle e,+|$ with $|s,+\rangle=\frac{1}{\sqrt{N}}\sum_{j}|s,j\rangle$ for $s=e,g$ [80]. The state $\rho^{+}$ has a large $l^{1}$-coherence $C_{l^{1}}(\rho^{+})=O(N)$, defined as $C_{l^{1}}(\rho)=\sum_{s,j,s^{\prime},j^{\prime}|(s,j)\neq(s^{\prime},j^{\prime})}|\langle s,j|\rho|s^{\prime},j^{\prime}\rangle|$, to which Ref. [57] attributed the dissipationless current. Now, note that the basis $\{|g,j\rangle,|e,j\rangle\}_{j=1}^{N}$, defining the $l^{1}$-coherence, is classical since there are only two jumps ($k=\pm$) regardless of $N$ and, for example, $\langle e,j^{\prime}|L_{+}|g,j\rangle=\sqrt{\gamma_{+}}=O(1)$.
@@
The hierarchy between conditions is exemplified by a state $\rho^{-}$ that has as much coherence regarding the classical eigenbasis as $\rho^{+}$ but does not yield dissipationless current (see Fig. 2(a)). Such a state is constructed by replacing $|s,+\rangle$ in the definition of $\rho^{+}$ with $|s,-\rangle=\frac{1}{\sqrt{N}}\sum_{j}(-1)^{j}|s,j\rangle$. For this state, the integrated fluxes and average escape rate read $\mathcal{T}_{s^{\prime}s}(\rho^{-})=\gamma_{s}p_{s}\chi(N)$ for $s\neq s^{\prime}$, $\mathcal{T}_{ss}(\rho^{-})=\frac{\gamma_{s}p_{s}}{2}(N-\chi(N))$, and $\bar{\lambda}(\rho^{-})=\frac{\gamma_{+}p_{g}+\gamma_{-}p_{e}}{2}(N+\chi(N))$, where $\gamma_{g}=\gamma_{+}$, $\gamma_{e}=\gamma_{-}$, and $\chi(N)$ is the remainder when $N$ is divided by $2$ [80]. Thus, $\rho^{-}$ does not satisfy either (Q1) and (Q2), and leads to $m_{H}(\rho^{-})=O(1)$. This example illustrates that the non-classicality represented via the quasiprobability is more restrictive than requiring a large amount of coherence, although both are just necessary conditions.
Finally, we remark that if we consider a non-classical basis including $|g,+\rangle$ and $|e,+\rangle$ (note that $\langle e,+|L_{+}|g,+\rangle=\sqrt{\gamma_{+}}N$), the state $\rho^{+}$ has no coherence regarding this basis. This fact shows that $m_{H}(\rho^{+})$ may exhibit anomalous scaling (that is, we may have dissipationless current) even if $\rho^{+}$ has no coherence if the referred basis is not classical (see Fig. 2(b)). Nonetheless, we see that $\rho^{+}$ fulfills either (Q1) or (Q2) for any eigenbasis, from the general discussion.
@@
Discussion.— We have explored fundamental connections between quantum thermodynamics and quasiprobability; in particular, we have derived a quantum TUR evaluating fluctuations through the TMH quasiprobability. We have also demonstrated that when discussing anomalous scaling of fluctuations in the TUR, which brings reduction of dissipation beyond classical limitations, non-classical behaviors of quasiprobability are more essential than quantum coherence. Although our results are derived based on the GKSL equation, which requires several conditions such as weak coupling between the system and the bath [61], the quasiprobability can be defined regardless of the dynamics; hence, we expect that the results can be insightful beyond the GKSL regime.
@@
To the best of our knowledge, our study is the first to bridge quasiprobabilities and the thermodynamic trade-off relations (except for a recent attempt [82] in a restricted situation, where negativity never appears). In addition to the TUR, several fundamental relations have been unveiled in the field, such as thermodynamic speed limits [83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93], bounds on asymmetry [94, 95, 96, 97], and fluctuation-response inequalities [98, 99, 100]. It would be intriguing to study how quasiprobabilities will help explore non-classical features in their quantum extensions.
@@ Acknowledgements.
Acknowledgments.— The authors are grateful to Hiroyasu Tajima and Ken Funo for fruitful discussions. K.Y. is supported by the Special Postdoctoral Researchers Program at RIKEN and JSPS KAKENHI Grant No. 24H00834. R.H. is supported by JSPS KAKENHI Grant No. JP24K16982. K.Y. and R.H. are supported by JST ERATO Grant No. JPMJER2302, Japan.
@@
$\mathbb{J}_{k}(\rho)$ $\coloneqq\begin{pmatrix}O&J_{-k}(\rho)\\ J_{k}(\rho)&O\end{pmatrix}$, (S1)
$\mathbb{F}_{k}(\rho)$ $\coloneqq\begin{pmatrix}O&F_{-k}(\rho)\\ F_{k}(\rho)&O\end{pmatrix}$, (S2)
@@
$J_{k}(\rho)$ $\coloneqq\frac{1}{2}(\gamma_{k}\tilde{L}_{k}\rho-\gamma_{-k}\rho\tilde{L}_{k}),$ (S3)
$F_{k}(\rho)$ $\coloneqq s_{k}\tilde{L}_{k}+[\tilde{L}_{k},\ln\rho].$ (S4)
@@
$\mathbb{J}(\rho)\coloneqq\bigoplus_{k}\mathbb{J}_{k}(\rho),\quad\mathbb{F}(\rho)\coloneqq\bigoplus_{k}\mathbb{F}_{k}(\rho),$ (S5)
@@
$\dot{\Sigma}(\rho)=\langle\mathbb{J}(\rho),\mathbb{F}(\rho)\rangle,$ (S6)
@@
$\nabla_{\mathbb{L}}A\coloneqq[I_{\mathfrak{h}}\otimes A,\mathbb{L}],$ (S7)
@@
$\mathbb{L}\coloneqq\bigoplus_{k}\begin{pmatrix}O&\tilde{L}_{-k}\\ \tilde{L}_{k}&O\end{pmatrix}$. (S8)
@@
$\Gamma=\bigoplus_{k}\begin{pmatrix}\gamma_{k}/2&0\\ 0&\gamma_{-k}/2\end{pmatrix}$. (S10)
@@
$\dot{\Sigma}(\rho)=\lVert\mathbb{F}(\rho)\rVert_{\Gamma\otimes\rho}^{2}.$ (S11)
@@
$\dot{\Sigma}(\rho)\lVert\nabla_{\mathbb{L}}X\rVert_{\Gamma\otimes\rho}^{2}\geq|\langle\mathbb{F}(\rho),\nabla_{\mathbb{L}}X\rangle_{\Gamma\otimes\rho}|^{2}.$ (S12)
@@
Appendix B Computational details of Example
We provide the details of the computation in the example in the main text. We aim to compute $\mathcal{T}_{s^{\prime}s}(\rho)$ and $\bar{\lambda}(\rho)$ for $\rho=\rho^{\pm}$ and show $m_{H}(\rho^{+})=O(N^{2})$ and $m_{H}(\rho^{-})=O(1)$. Further, we also compute the current $J_{H}^{\mathrm{d}}(\rho^{-})$ and the EPR $\dot{\Sigma}(\rho^{-})$ and confirm the inequality. Here, the model we consider consists of Hamiltonian $H=\omega\sum_{j}|e,j\rangle\langle e,j|$ and jump operators $L_{+}=\sqrt{\gamma_{+}}\sum_{j,j^{\prime}}|e,j\rangle\langle g,j^{\prime}|$ and $L_{-}=\sqrt{\gamma_{-}}\sum_{j,j^{\prime}}|g,j\rangle\langle e,j^{\prime}|$. The states of interest $\rho^{\pm}$ are generated as $\rho^{\pm}=\sum_{s=g,e}p_{s}|s,\pm\rangle\langle s,\pm|$, where $|s,\pm\rangle=\frac{1}{\sqrt{N}}\sum_{j}(\pm 1)^{j}|s,j\rangle$ and $p_{s}$ satisfies $p_{s}\geq 0$ and $p_{g}+p_{e}=1$.
@@ Appendix B Computational details of Example
$T_{s^{\prime}j^{\prime}sj}(\rho)=\frac{1}{2}\mathop{\mathrm{tr}}\nolimits\big(\{\mathcal{D}^{\dagger}\Pi_{s^{\prime}j^{\prime}},\Pi_{sj}\}\rho\big).$ (S13)
@@
$\sigma_{s,j,j^{\prime}}$ $\coloneqq\frac{1}{2}\big(|s,j\rangle\langle s,j^{\prime}|+|s,j^{\prime}\rangle\langle s,j|\big),$ (S14)
$\Lambda_{s,j}$ $\coloneqq\sum_{j^{\prime}}\sigma_{s,j,j^{\prime}}.$ (S15)
@@
$\sigma_{s,j,j^{\prime}}=\sigma_{s,j^{\prime},j},\;\;\sigma_{s,j,j}=\Pi_{s,j},$ (S16)
$\mathop{\mathrm{tr}}\nolimits(\sigma_{s,j,j^{\prime}}\rho)=\mathrm{Re}\langle s,j|\rho|s,j^{\prime}\rangle,$ (S17)
$\sum_{j}\Lambda_{s,j}=\sum_{j,j^{\prime}}|s,j\rangle\langle s,j^{\prime}|,$ (S18)
$\frac{1}{2}\sum_{j^{\prime}}\{\Lambda_{s^{\prime},j^{\prime}},\Pi_{s,j}\}=\delta_{s,s^{\prime}}\Lambda_{s,j}.$ (S19)
@@
$L_{+}^{\dagger}\Pi_{s,j}L_{+}$ $=\gamma_{+}\delta_{s,e}\sum_{j^{\prime}}\Lambda_{g,j^{\prime}},$ (S20)
$L_{-}^{\dagger}\Pi_{s,j}L_{-}$ $=\gamma_{-}\delta_{s,g}\sum_{j^{\prime}}\Lambda_{e,j^{\prime}},$ (S21)
@@
$L_{+}^{\dagger}L_{+}$ $=\gamma_{+}N\sum_{j^{\prime}}\Lambda_{g,j^{\prime}},$ (S22)
$L_{-}^{\dagger}L_{-}$ $=\gamma_{-}N\sum_{j^{\prime}}\Lambda_{e,j^{\prime}}.$ (S23)
@@
$+\delta_{s,g}\bigg(\gamma_{-}\sum_{j^{\prime}}\Lambda_{e,j^{\prime}}-\gamma_{+}N\Lambda_{g,j}\bigg).$ (S24)
@@
$+\delta_{s^{\prime},g}\delta_{s,e}\gamma_{-}\Lambda_{e,j}-\frac{\delta_{s^{\prime},g}\delta_{s,g}}{2}\gamma_{+}N(\delta_{j,j^{\prime}}\Lambda_{g,j}+\sigma_{g,j,j^{\prime}}).$ (S25)
@@
$T_{ej^{\prime}gj}(\rho)$ $=\gamma_{+}\sum_{j^{\prime\prime}}\mathrm{Re}\langle g,j|\rho|g,j^{\prime\prime}\rangle$ (S26)
$T_{gj^{\prime}ej}(\rho)$ $=\gamma_{-}\sum_{j^{\prime\prime}}\mathrm{Re}\langle e,j|\rho|e,j^{\prime\prime}\rangle$ (S27)
@@
$\phantom{===}+\delta_{j,j^{\prime}}\sum_{j^{\prime\prime}}\mathrm{Re}\langle g,j|\rho|g,j^{\prime\prime}\rangle\bigg)$ (S28)
@@
$\phantom{===}+\delta_{j,j^{\prime}}\sum_{j^{\prime\prime}}\mathrm{Re}\langle e,j|\rho|e,j^{\prime\prime}\rangle\bigg).$ (S29)
@@
$T_{ej^{\prime}gj}(\rho^{+})=p_{g}\gamma_{+},\;\;T_{ej^{\prime}gj}(\rho^{-})=(-1)^{j+1}p_{g}\gamma_{+}\frac{\chi(N)}{N},$ (S30)
@@
$\mathcal{T}_{eg}(\rho^{+})=p_{g}\gamma_{+}N^{2},\quad\mathcal{T}_{eg}(\rho^{-})=p_{g}\gamma_{+}\chi(N).$ (S31)
@@
$T_{gj^{\prime}gj}(\rho^{+})$ $=-\frac{\gamma_{+}p_{g}}{2}(1+\delta_{j,j^{\prime}}N)$ (S32)
$T_{gj^{\prime}gj}(\rho^{-})$ $=-\frac{(-1)^{j}\gamma_{+}p_{g}}{2}((-1)^{j^{\prime}}-\delta_{j,j^{\prime}}\chi(N)).$ (S33)
@@
$\mathcal{T}_{gg}(\rho^{+})$ $=-\frac{\gamma_{+}p_{g}}{2}N(N-1),$ (S34)
$\mathcal{T}_{gg}(\rho^{-})$ $=\frac{\gamma_{+}p_{g}}{2}(N-\chi(N))$ (S35)
@@
$\sum_{j,j^{\prime}(\neq j)}(-1)^{j}(-1)^{j^{\prime}}$ $=\bigg(\sum_{j}(-1)^{j}\bigg)^{2}-\sum_{j}(-1)^{2j}$ (S36)
$=\chi(N)-N.$ (S37)